A line of length 10 and one end is at point (2,-3);if the abscissa of the other end be 10,prove that it's ordinate must be 3 or -9

Grade:12th pass

1 Answers

abdul alim
6 years ago
aplly distance formula let the ordinae of other point be y then the point will become (10,y) and one point we have (2,-3)(10-2)^2+(y+3)^2=10^264+y^2+9+6y=100y^2+6y-27=0(y-3)(y+9)=0thus y=3 and -9 hence peoved